OpenStudy (anonymous):

Which is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer? 6.2x + 1.2 = 8.9 A. 1.3 B. 1.4 C. 1.6 D. 1.7

OpenStudy (anonymous):

Do you know how to solve an equation algebraically? Subtract the 1.2 to the other side, so you get 6.2x = 9.7. Divide each side by 6.2 and round.

OpenStudy (anonymous):

Oh, my bad - round the numbers first (0-4, number before the decimal, 5-9, number above before decimal. So, 6.2 = 6, and 8.9 = 10, etc.).

OpenStudy (anonymous):

But then use the same solving process.

OpenStudy (texaschic101):

8.9 = 9

OpenStudy (anonymous):

"Which is the best ESTIMATE of x BASED on rounding the constants and coefficients in the equation to the nearest integer?" 6.2x + 1.2 = 8.9 becomes 6x + 1 = 9 isolate x term 6x = 9-1 isolate x x = 8/6

OpenStudy (texaschic101):

6x + 1 = 9 6x = 8 x = 8/6 rounds to 4/3 = 1 1/3 or 1.3